# Meta-GGAs vs. Hybrid Functionals for Point Defects: The Best of Both Worlds Applied to Layered MnO2, NiO2 and KCoO2

### v2

@@ Abstract
Defects in a material can significantly tune properties and enhance utility. Hybrid functionals like HSE06 are often used to describe solids with such defects. However, geometry optimization (including accounting for effects such as Jahn-Teller distortion) using hybrid functionals is challenging for the large supercells needed for defect study. The proposed r2SCAN+rVV10+U+$\mathrm{U_{d}}$ method, which is computationally much cheaper and faster than hybrid functionals, can successfully describe defects in materials with the proper choice of U (for the d orbitals of the host atom) and $\mathrm{U_{d}}$ (for those of the defect atom), as shown here for small polaron defects in layered transition-metal oxides. We use a literature value of U or $\mathrm{U_{d}}$ appropriate to a given transition-metal ion and its oxidation state. The layered materials birnessite ($\mathrm{K_{n}MnO_{2}},n=0.03$) and $\mathrm{K_{n}NiO_{2}},n=0.03$, with one K atom intercalated between layers in a supercell, are found to have one localized occupied $\mathrm{e_{g}}$ polaronic state on the transition metal ion reduced by the insertion of the K atom, when the geometry is calculated as above using published U values. The expected Jahn-Teller distortion is not observed when U=$\mathrm{U_{d}}$=0. Layered cobalt oxide with additional potassium ions intercalated ($\mathrm{K_{n}CoO_{2}},n=1.03$) is different, due to a dramatic difference in electronic configuration of the defected Co(II) ion: A single extra K atom in the supercell leads to four localized electrons in the band gap, using standard U values, and even for U=$\mathrm{U_{d}}$=0.
@@ I Introduction
This work is part of a larger project on the computational identification of alkali-atom-intercalated layered materials as promising catalysts for the oxygen evolution reaction (OER) of water splitting for clean hydrogen production. Earlier work Peng et al. (2017); Ding et al. (2021) showed the importance of small-polaron defects formed by transfer of an electron from an intercalated alkali atom to a neighboring manganese ion. Calculation of the polaron often requires a nonlocal density functional for the exchange-correlation energy, and geometry optimization in a large supercell. We found that hybrid functionals as used in Ref. Peng et al. (2017) are too expensive for a broad materials search. In this article, we show that an alternative r2SCAN+rVV10+U+$\mathrm{U_{d}}$ approach can achieve comparable accuracy at much lower cost, and may be useful for other point defects in a range of materials. We also found that, while one intercalated K atom creates one localized electron in the energy gap of layered MnO2 and NiO2, it can create four in layered KCoO2. We further found that, while r2SCAN needs a +U correction to create a polaron in MnO2 and NiO2, it does not need one in KCoO2. Our broader computational and experimental materials search is now underway.
@@
Hybrid functionals, which combine a fraction of exact exchange, such as HSE06 Krukau et al. (2006a) which utilizes 25% exact exchange and 75% PBE exchange in the short range along with full PBE exchange in the long range, experience less SIE and often provide a more accurate description of the electronic structure of materials. Hybrid functionals have been successfully employed in studying defects in solids. For instance, Peng et al. Peng et al. (2017) effectively investigated polaron-like defects in birnessite (Fig. 1, $\mathrm{K_{n}MnO_{2}},n<1$). Nevertheless, the inclusion of exact exchange in hybrid functionals renders them computationally expensive. The structural relaxation using hybrid functionals becomes particularly costly for a reasonably sized supercell with localized defects, with computational expenses rapidly escalating as the supercell size increases toward better simulation of defects. Consequently, the structural optimization of large supercells becomes impractical. Another challenge associated with hybrid functionals is that a material-independent exact exchange mixing parameter is not determined through any exact condition, nor is the range-separation parameter in range-separated hybrids Sun et al. (2016). Peng et al. determined a mixing parameter of 0.22 to study defects in birnessite using HSE06 Peng et al. (2017), deviating from the original value of 0.25. Additionally, determining a mixing parameter for semiconductors that may not be suitable to metals or insulators Henderson et al. (2011) adds to the challenge of finding an appropriate parameter for the system under study when employing hybrid functionals.
@@
$\mathrm{r^{2}SCAN}$ Furness et al. (2020a) is a recently developed meta-generalized gradient approximation that reinstates exact constraint adherence to rSCAN Bartók and Yates (2019), preserving the numerical efficiency of rSCAN while simultaneously restoring the transferable accuracy of SCAN Furness et al. (2020b). Several studies have demonstrated that SCAN predicts geometries and other properties as well as or even better than hybrid functionals. Sun et al. Sun et al. (2016) showed that SCAN accurately predicts geometries and energies of diversely bonded materials and molecules, matching or surpassing the accuracy of computationally expensive hybrid functionals. Another study by Saÿnick and Cocchi Saÿnick and Cocchi (2021) on cesium-based photocathode materials $\mathrm{Cs_{3}Sb}$ and $\mathrm{Cs_{2}Te}$ reported excellent performance of SCAN and HSE06 for both structural and electronic properties. A recent paper on the arXiv Wickramaratne and Lyons reports that, while SCAN may not reliably describe the properties of deep defects and small polarons in several semiconductors and insulators, it yields remarkably good agreement with experimental structural parameters for materials like ZnO, GaN, $\mathrm{Ga_{2}O_{3}}$ and NaF. Additionally, a study by Varadwaj and Miyake Varadwaj and Miyake (2022) on the geometrical, electronic, and optical properties of vanadium dioxide found that SCAN and SCAN-rVV10 can adequately predict the most important geometrical and optoelectronic properties of $\mathrm{VO_{2}}$. Numerous related studies further support the idea that SCAN successfully describes the structure and other properties of materials. Given that $\mathrm{r^{2}SCAN}$ closely agrees with SCAN in accuracy, we expect that $\mathrm{r^{2}SCAN}$ would exhibit similar accuracy in the aforementioned studies.
@@
By construction, $\mathrm{r^{2}SCAN}$ can exhibit very small or negligible SIE (as reflected by its smaller Hubbard U correction) and performs at the level of hybrid functionals but demands less computational resource and time. In this study, we demonstrate that the $\mathrm{r^{2}SCAN+rVV10+U+U_{d}}$ functional can effectively describe the defects in materials at the level of the hybrid functional HSE06, or possibly even better. Here, U represents the Hubbard U correction of Anisimov and collaborators Anisimov et al. (1991, 1993); Solovyev et al. (1994), applied to transition metal sites other than the defect site, while $\mathrm{U_{d}}$ is the correction applied to the defect site. Cococcioni and de Gironcoli Cococcioni and de Gironcoli (2005) showed that the +U correction can be regarded as a many-electron self-interaction correction that, like the PPLB condition Perdew et al. (1982), penalizes non-integer electron number on a localized orbital to which it is applied.
@@
However, this situation is not universal, as we can see that U values for Mn for $\mathrm{r^{2}SCAN}$ are 1.5 eV, 2.1 eV and 1.8 eV for Mn ions with average oxidation states +3.5, +2.5 and +2.33 respectively Swathilakshmi et al. (2023). Here we see that Mn in the +2.5 average OS needs more U correction than Mn in the +2.33 average OS, which does not follow the trend observed for vanadium, although U for Mn in the +3.5 average OS supports the trend within the same elemental series. However it is obvious for the Mn oxide system that Mn ions in different oxidation states need different U corrections. Another example for Co is found in Ref. Long et al. (2020), which found U=3.0 eV as a correction to SCAN for the oxidation reaction 6CoO+$\mathrm{O_{2}\rightarrow 2Co_{3}O_{4}}$ (in which the Co ions are in the +2 and +2.67 oxidation states), but also found that SCAN with this U wrongly predicted the meta-stable crystal O1-CoO2 (in which the Co oxidation state is +4) to be non-metallic, while SCAN with U=0 correctly predicted it to be a metal. Thus the U value can vary with the oxidation state of ions in the system. This indicates that if a system has transition metal ions in different OS, the proper way to describe such a system would be to apply different U corrections to the transition metal ions based on their OS. This could be because DFAs make different SIE for ions in different OS.
@@ II.1 Layered MnO2
Figure 3: r2SCAN+rVV10+U+$\mathrm{U_{d}}$ spin-resolved density of states per atom in (A) pristine MnO2 with U=$\mathrm{U_{d}}$=1.8 eV and in a single K-intercalated MnO2 with (B) U = $\mathrm{U_{d}}$ = 0.0 eV, (C) U= $\mathrm{U_{d}}$ =1.8 eV, (D) U=1.71 eV and $\mathrm{U_{d}}$ =1.58 eV. The geometry has been optimized in r2SCAN+rVV10+U+$\mathrm{U_{d}}$, which makes the interlayer spacing  5.26 Å. The insets in Figs. A and B show the DOS near the 0 eV region. The inset of Fig. B shows that for U=$\mathrm{U_{d}}$=0 the extra electron from the intercalated K goes to the bottom of the conduction band, making a semi-metal.
@@
Motivated by studies indicating that U depends on OS, and that generally U decreases with an increase in OS, we determined U for the r2SCAN+rVV10+U functional for Mn(III) and Mn(IV) by comparing r2SCAN+rVV10+U’s magnetic moment with the HSE06+D3 magnetic moment. For this we used pristine KMnO2 and MnO2 systems. First, we relaxed the structure using HSE06+D3 and used the structure to get the magnetic moment due to spin charge density within the Wigner-Seitz sphere around an Mn ion. We used the HSE06+D3 geometry for r2SCAN+rVV10+U and slowly increased U in small steps until we obtained the same magnetic moment as HSE06+D3. This method gave U values of 1.71 eV and 1.58 eV for Mn(III) and Mn(IV) ions, respectively. These values are close to the values in Ref. Swathilakshmi et al. (2023).
@@
We also performed HSE06+D3 calculations on the r2SCAN+rVV10+U+$\mathrm{U_{d}}$ geometries, and the resulting DOS are plotted in Fig. 4. The D3 dispersion correction was used for the HSE06 functional with new D3 damping parameters a1 = 0.383, a2 = 5.685, and s8 = 2.310 generated for the HSE06 functional Jonas and Grimme (2014). For HSE06+D3 calculations, we used an exact exchange mixing parameter ($\alpha$) of 0.25. This is because we found that the r2SCAN+rVV10+U geometry in much better agreement with the HSE06+D3 geometry for $\alpha$=0.25 than for 0.22. More details are in the Supplementary Material.
@@
r2SCAN+rVV10+U+$\mathrm{U_{d}}$ with U=$\mathrm{U_{d}}$=1.8 eV, and U=1.71 eV and $\mathrm{U_{d}}$=1.58 eV give similar DOS. HSE06+D3 DOS are also similar for those geometries, with the polaron appearing deep in the gap.
@@ II.2 Layered NiO2
We began the DOS calculation using the r2SCAN+rVV10+U+$\mathrm{U_{d}}$ functional for both pristine and K-intercalated NiO2 systems. The r2SCAN U value for Ni ions was recently determined by Swathilakshmi et al. Swathilakshmi et al. (2023). For the pristine NiO2, r2SCAN+rVV10+U+$\mathrm{U_{d}}$ with U=$\mathrm{U_{d}}$=2.1 eV, correctly predicts a non-magnetic ground state with a band gap of 1.62 eV, as shown in Fig. 6A. The d-states of Ni(IV) in pristine NiO2 are split into t2g and eg states, with t2g states completely occupied and eg states empty, as shown in Fig. 8A. As in the birnessite case, we expect that adding an extra K atom between layers would transfer an electron from the inserted K atom to a Ni site, forming a defect. The defected Ni site would undergo Jahn-Teller distortion, localizing the electron in the eg state, as shown in Fig. 8B.
Figure 6: r2SCAN+rVV10+U+$\mathrm{U_{d}}$ spin-resolved density of states per atom in (A) pristine NiO2 with Ni(IV) ions with U=$\mathrm{U_{d}}$=2.1 eV and in a single K-intercalated NiO2 with (B) U=$\mathrm{U_{d}}$ =0.0 eV, (C) U=$\mathrm{U_{d}}$=2.1 eV, (D) U= $\mathrm{U_{d}}$ =2.06 eV. The geometry was optimized in r2SCAN+rVV10+U+$\mathrm{U_{d}}$, which makes the interlayer spacing   5.42 Å. Here we can see one localized occupied state and several localized unoccupied states on the defect Ni(III) cation. The insets in Figs. A and B show the DOS near the 0 eV region. The inset in Fig. B shows that for U=$\mathrm{U_{d}}$=0 the extra electron from the intercalated K goes to the bottom of the conduction band, making a semi-metal.
@@
For a K-intercalated NiO2, we initially used U=$\mathrm{U_{d}}$=0.0 eV to observe how r2SCAN+rVV10 without U correction performs for this system, and we obtained the DOS as plotted in Fig. 6B. Again, this choice of U and $\mathrm{U_{d}}$ failed to describe the defect state, as can be seen from the inset in Fig. 6B. The extra electron goes to a delocalized state at the bottom of the conduction band, with fractional occupation on each Ni ion in the supercell, and no individual geometric distortion is observed at any nickel atom. We then used the available U value by setting U=$\mathrm{U_{d}}$=2.1 eV and obtained the DOS, as plotted in Fig. 6C. This choice of U values forms a defect at one of the Ni sites, evidenced by the expected geometric distortion (see Fig. S3 B). The polaron is localized and appears in the gap just below the CB minimum. The formation of the polaron at the VB edge indicates that a K -intercalated NiO2 could show OER catalytic activity similar to that of birnessite by lowering the overpotential. Most of the charge from the K atom is transferred to this defect site, which has a magnetic moment of 0.786 $\mu$B . However, we observe a Ni site in another layer picking up a small but nonzero magnetic moment of 0.137 $\mu$B, with the magnetic moment of all remaining Ni sites smaller than 0.03 $\mu$B. This suggests that we might need different values of U and $\mathrm{U_{d}}$ for r2SCAN+rVV10+U+$\mathrm{U_{d}}$ to accurately describe charge transfer to the defect site in a K-intercalated NiO2 system.
[…]
We also determined the U value of 2.06 eV for the Ni(III) ion by equating the HSE06+D3 magnetic moment to the r2SCAN+rVV10+U magnetic moment, as discussed in the Supplementary Material. As this value is not very different from 2.1 eV, we got a similar r2SCAN+rVV10+U+$\mathrm{U_{d}}$ result with U=$\mathrm{U_{d}}$=2.06 eV compared to U=$\mathrm{U_{d}}$=2.1 eV. Fig. 6D shows r2SCAN+rVV10+U+$\mathrm{U_{d}}$ DOS with U=$\mathrm{U_{d}}$=2.06 eV.

### v4

# Comparing Meta-GGAs, +U Corrections, and Hybrid Functionals for Polaronic Point Defects in Layered MnO2, NiO2, and KCoO2

@@ Abstract
Defects in a material can significantly tune properties and enhance utility. Hybrid functionals like HSE06 are often used to describe solids with such defects. However, geometry optimization (including accounting for effects such as Jahn-Teller distortion) using hybrid functionals is challenging for the large supercells needed for defect study. The proposed r2SCAN+rVV10+U+$\mathrm{U_{d}}$ method, which is computationally much cheaper and faster than hybrid functionals, can successfully describe defects in materials with the proper choice of U (for the d orbitals of the host atom) and $\mathrm{U_{d}}$ (for those of the defect atom), as shown here for small polaron defects in layered transition-metal oxides. For a range of U and $\mathrm{U_{d}}$ around literature values (from solid-state reaction energies) for a given transition-metal ion and its oxidation state, we find that this approach predicts localized polaronic states in band gaps, as hybrid functionals do. The layered materials birnessite ($\mathrm{K_{n}MnO_{2}},n=0.03$) and $\mathrm{K_{n}NiO_{2}},n=0.03$, with one K atom intercalated between layers in a supercell, are found to have one localized occupied $\mathrm{e_{g}}$ polaronic state on the transition metal ion reduced by the insertion of the K atom, when the geometry is calculated as above using published U values. The expected Jahn-Teller distortion is not observed when U=$\mathrm{U_{d}}$=0. Layered cobalt oxide with additional potassium ions intercalated ($\mathrm{K_{n}CoO_{2}},n=1.03$) is different, due to a dramatic difference in electronic configuration of the defected Co(II) ion: A single extra K atom in the supercell leads to four localized electrons in the band gap, using standard U values, and even for U=$\mathrm{U_{d}}$=0.
@@ I Introduction
This work is part of a larger project on the computational identification of alkali-atom-intercalated layered materials as promising catalysts for the oxygen evolution reaction (OER) of water splitting for clean hydrogen production. Earlier work Peng et al. (2017); Ding et al. (2021) showed the importance of small-polaron defects formed by transfer of an electron from an intercalated alkali atom to a neighboring manganese ion. Electron flow in catalysis seems to be facilitated by a localized polaronic state in the energy gap close to the bottom of the conduction band. Calculation of the polaron often requires a nonlocal density functional for the exchange-correlation energy, and geometry optimization in a large supercell. We found that hybrid functionals as used in Ref. Peng et al. (2017) are too expensive for a broad materials search. In this article, we show that an alternative r2SCAN+rVV10+U+$\mathrm{U_{d}}$ approach can predict localized polaronic states at much lower cost, and may be useful for other point defects in a range of materials. We also found that, while one intercalated K atom creates one localized electron in the energy gap of layered MnO2 and NiO2, it can create four in layered KCoO2. We further found that, while r2SCAN needs a +U correction to create a polaron in MnO2 and NiO2, it does not need one in KCoO2. Our broader computational and experimental materials search is now underway.
@@
Hybrid functionals, which combine a fraction of exact exchange, such as HSE06 Krukau et al. (2006) which utilizes 25% exact exchange and 75% PBE exchange in the short range along with full PBE exchange in the long range, experience less SIE and often provide a more accurate description of the electronic structure of materials. Hybrid functionals have been successfully employed in studying defects in solids. For instance, Peng et al. Peng et al. (2017) effectively investigated polaron-like defects in birnessite (Fig. 1, $\mathrm{K_{n}MnO_{2}},n<1$). Nevertheless, the inclusion of exact exchange in hybrid functionals renders them computationally expensive. The structural relaxation using hybrid functionals becomes particularly costly for a reasonably sized supercell with localized defects, with computational expenses rapidly escalating as the supercell size increases toward better simulation of defects. Consequently, the structural optimization of large supercells becomes impractical. Another challenge associated with hybrid functionals is that a material-independent exact exchange mixing parameter is not determined through any exact condition, nor is the range-separation parameter in range-separated hybrids Sun et al. (2016). Peng et al. determined a mixing parameter of 0.22 to study defects in birnessite using HSE06 Peng et al. (2017), deviating from the original value of 0.25. Additionally, determining a mixing parameter for semiconductors that may not be suitable to metals or insulators Henderson et al. (2011) adds to the challenge of finding an appropriate parameter for the system under study when employing hybrid functionals.
@@
$\mathrm{r^{2}SCAN}$ Furness et al. (2020) is a recently developed meta-generalized gradient approximation that reinstates exact constraint adherence to rSCAN Bartók and Yates (2019), preserving the numerical efficiency of rSCAN while simultaneously restoring the transferable accuracy of SCAN Sun et al. (2015). Several studies have demonstrated that SCAN predicts geometries and other properties as well as or even better than hybrid functionals. Sun et al. Sun et al. (2016) showed that SCAN accurately predicts geometries and energies of diversely bonded materials and molecules, matching or surpassing the accuracy of computationally expensive hybrid functionals. Another study by Saÿnick and Cocchi Saÿnick and Cocchi (2021) on cesium-based photocathode materials $\mathrm{Cs_{3}Sb}$ and $\mathrm{Cs_{2}Te}$ reported excellent performance of SCAN and HSE06 for both structural and electronic properties. A recent paper on the arXiv Wickramaratne and Lyons reports that, while SCAN may not reliably describe the properties of deep defects and small polarons in several semiconductors and insulators, it yields remarkably good agreement with experimental structural parameters for materials like ZnO, GaN, $\mathrm{Ga_{2}O_{3}}$ and NaF. Additionally, a study by Varadwaj and Miyake Varadwaj and Miyake (2022) on the geometrical, electronic, and optical properties of vanadium dioxide found that SCAN and SCAN-rVV10 can adequately predict the most important geometrical and optoelectronic properties of $\mathrm{VO_{2}}$. Numerous related studies further support the idea that SCAN successfully describes the structure and other properties of materials. Given that $\mathrm{r^{2}SCAN}$ closely agrees with SCAN in accuracy, we expect that $\mathrm{r^{2}SCAN}$ would exhibit similar accuracy in the aforementioned studies.
@@
By construction, $\mathrm{r^{2}SCAN}$ can exhibit considerably less self-interaction error (as reflected in its smaller U value Gautam and Carter (2018)) than standard GGAs. (The r2SCAN effective one-electron potential can still show substantial self-interaction error.) In this study, we demonstrate that the $\mathrm{r^{2}SCAN+rVV10+U+U_{d}}$ functional can predict the existence of localized polaronic states at defects, as the HSE06 hybrid functional does. Here, U represents the Hubbard U correction of Anisimov and collaborators Anisimov et al. (1991, 1993); Solovyev et al. (1994), applied to transition metal sites other than the defect site, while $\mathrm{U_{d}}$ is the correction applied to the defect site. Cococcioni and de Gironcoli Cococcioni and de Gironcoli (2005) showed that the +U correction can be regarded as a many-electron self-interaction correction that, like the PPLB condition Perdew et al. (1982), penalizes non-integer electron number on a localized orbital to which it is applied.
@@
[…]
@@
From a chemical perspective Atkins et al. , for electronic configurations (3d)n $(4\leq n\leq 7)$, the spin state of a transition metal ion M in an MO6 motif arises from a competition between crystal field effects which favor the low-spin state and internal effects which favor the high-spin state. The crystal field effect (which in the absence of spin removes the degeneracy of three t2g and two eg orbitals) depends on the lattice geometry, including any Jahn-Teller distortion of the MO6 octahedron. The t2g orbitals have lobes between the bond axes, minimizing the repulsion between the metal ion and its surrounding oxygen ions. The ions we consider here are in oxidation states Mn(IV) (d3), Mn(III) (d4), Ni(IV) (d6), Ni(III) (d7), Co(III) (d6), and Co(II) (d7).
@@ II.1 Layered MnO2
Figure 3: r2SCAN+rVV10+U+$\mathrm{U_{d}}$ spin-resolved density of states per atom in (A) pristine MnO2 with U=$\mathrm{U_{d}}$=1.8 eV and in a single K-intercalated MnO2 with (B) U = $\mathrm{U_{d}}$ = 0.0 eV, (C) U= $\mathrm{U_{d}}$ =1.8 eV, (D) U=1.71 eV and $\mathrm{U_{d}}$ =1.58 eV. The geometry has been optimized in r2SCAN+rVV10+U+$\mathrm{U_{d}}$, which makes the interlayer spacing $\approx$
5.26 Å. The insets in Figs. A and B show the DOS near the 0 eV region. The inset of Fig. B shows that for U=$\mathrm{U_{d}}$=0 the extra electron from the intercalated K goes to the bottom of the conduction band, making a semi-metal.
@@
Motivated by studies indicating that U depends on OS, and that generally U decreases with an increase in OS, we determined U for the r2SCAN+rVV10+U functional for Mn(III) and Mn(IV) by matching r2SCAN+rVV10+U’s magnetic moment for a pristine material to the HSE06+D3 magnetic moment (Supplementary Material (SM) SM- ). (All magnetic moments reported in this article arise from spin density inside the Wigner-Seitz sphere.) This method gave U values of 1.71 eV and 1.58 eV for Mn(III) and Mn(IV) ions, respectively. These values are close to the values in Ref. Swathilakshmi et al. (2023).
@@
We also performed HSE06+D3 calculations on the r2SCAN+rVV10+U+$\mathrm{U_{d}}$ geometries, and the resulting DOS are plotted in Fig. 4. The D3 dispersion correction was used for the HSE06 functional with new D3 damping parameters a1 = 0.383, a2 = 5.685, and s8 = 2.310 generated for the HSE06 functional Jonas and Grimme (2014). For HSE06+D3 calculations, we used an exact exchange mixing parameter ($\alpha$) of 0.25. This is because we found that the r2SCAN+rVV10+U geometry in much better agreement with the HSE06+D3 geometry for $\alpha$=0.25 than for 0.22. More details are in the SM SM- .
@@ II.2 Layered NiO2
We began the DOS calculation using the r2SCAN+rVV10+U+$\mathrm{U_{d}}$ functional for both pristine and K-intercalated NiO2 systems. The r2SCAN U value for Ni ions was recently determined by Swathilakshmi et al. Swathilakshmi et al. (2023). For the pristine NiO2, r2SCAN+rVV10+U+$\mathrm{U_{d}}$ with U=$\mathrm{U_{d}}$=2.1 eV, correctly predicts a non-magnetic ground state with a band gap of 1.62 eV, as shown in Fig. 6A. As in the birnessite case, we expect that adding an extra K atom between layers would transfer an electron from the inserted K atom to a Ni site, forming a defect. The defected Ni site would undergo Jahn-Teller distortion, localizing the electron in the eg state, as shown in Fig. 8B.
Figure 6: r2SCAN+rVV10+U+$\mathrm{U_{d}}$ spin-resolved density of states per atom in (A) pristine NiO2 with Ni(IV) ions with U=$\mathrm{U_{d}}$=2.1 eV and in a single K-intercalated NiO2 with (B) U=$\mathrm{U_{d}}$ =0.0 eV, (C) U=$\mathrm{U_{d}}$=2.1 eV, (D) U= $\mathrm{U_{d}}$ =2.06 eV. The geometry was optimized in r2SCAN+rVV10+U+$\mathrm{U_{d}}$, which makes the interlayer spacing $\approx$ 5.42 Å. Here we can see one localized occupied state and several localized unoccupied states on the defect Ni(III) cation. The insets in Figs. A and B show the DOS near the 0 eV region. The inset in Fig. B shows that for U=$\mathrm{U_{d}}$=0 the extra electron from the intercalated K goes to the bottom of the conduction band, making a semi-metal.
@@
For a K-intercalated NiO2, we initially used U=$\mathrm{U_{d}}$=0.0 eV to observe how r2SCAN+rVV10 without U correction performs for this system, and we obtained the DOS as plotted in Fig. 6B. Again, this choice of U and $\mathrm{U_{d}}$ failed to describe the defect state, as can be seen from the inset in Fig. 6B. The extra electron goes to a delocalized state at the bottom of the conduction band, with fractional occupation on each Ni ion in the supercell, and no individual geometric distortion is observed at any nickel atom. We then used the available U value by setting U=$\mathrm{U_{d}}$=2.1 eV and obtained the DOS, as plotted in Fig. 6C. This choice of U values forms a defect at one of the Ni sites, evidenced by the expected geometric distortion (see Fig. S3 B). The polaron is localized and appears in the gap just below the CB minimum. The formation of the polaron at the VB edge indicates that a K -intercalated NiO2 could show OER catalytic activity similar to that of birnessite by lowering the overpotential. Most of the charge from the K atom is transferred to this defect site, which has a magnetic moment of 0.786 $\mathrm{\mu_{B}}$ However, we observe a Ni site in another layer picking up a small but nonzero magnetic moment of 0.137 $\mathrm{\mu_{B}}$, with the magnetic moment of all remaining Ni sites smaller than 0.03 $\mathrm{\mu_{B}}$. This suggests that we might need different values of U and $\mathrm{U_{d}}$ for r2SCAN+rVV10+U+$\mathrm{U_{d}}$ to accurately describe charge transfer to the defect site in a K-intercalated NiO2 system.
The determination of U=2.1 eV for Ni for r2SCAN involves Ni in +2 and +3 states Swathilakshmi et al. (2023), with an average OS of +2.5. Since the precise U values for Ni(III) and Ni(IV) states are not available, we set $\mathrm{U_{d}}$=2.1 eV for NI(III) as the Ni (III) OS is close to +2.5 and searched for different U values for Ni(IV). We observed that increasing the U value beyond 2.3 eV transfers more and more charge to a Ni site in a layer different from that of the defect site. So, we lowered the U values (U=1.0 eV and 0.0 eV ) and found that lowering the U values slightly improves the solution by lowering the magnetic moment of a Ni site in another layer to 0.06 $\mathrm{\mu_{B}}$ without significant change in magnetic moment of the defect site. The DOS plot for U=0.0 eV and $\mathrm{U_{d}}$=2.1 eV do not look different from Fig. 6C. We also increased the $\mathrm{U_{d}}$ value, keeping U fixed at 2.1 eV, and observed an increase in the magnetic moment of the defect site, reaching 1 $\mathrm{\mu_{B}}$ for U=3.8 eV. These are the magnetic moments due to the spin density inside the Wigner-Seitz sphere. It is important to note that the magnetic moment of the defect site 1.0 does not guarantee the full transfer of an electron from the intercalated K atom to the defect site, as this is a magnetic moment inside the Wigner-Seitz sphere, which does not reflect the actual magnetic moment of the defect site. However, this analysis shows that one can adjust the U and $\mathrm{U_{d}}$ values for more charge transfer to the defect site where necessary.
We also determined the U value of 2.06 eV for the Ni(III) ion by equating the HSE06+D3 magnetic moment to the r2SCAN+rVV10+U magnetic moment, as discussed in the SM SM- . As this value is not very different from 2.1 eV, we got a similar r2SCAN+rVV10+U+$\mathrm{U_{d}}$ result with U=$\mathrm{U_{d}}$=2.06 eV compared to U=$\mathrm{U_{d}}$=2.1 eV. Fig. 6D shows r2SCAN+rVV10+U+$\mathrm{U_{d}}$ DOS with U=$\mathrm{U_{d}}$=2.06 eV.